OpenStudy (anonymous):

-n/5= -5 N = -25 N = -1 N = 25 ok wtf how do i do this do i have two negatives or one or wtfffffffff

OpenStudy (anonymous):

solve for N

OpenStudy (anonymous):

5-N/5 = 0 N = 25

OpenStudy (anonymous):

explain it u dippellete

OpenStudy (anonymous):

-N = (-5)(5) -N = -25 N = 25

OpenStudy (anonymous):

whoa computer freaked out sorry explain it

OpenStudy (anonymous):

-n/5=5 -n=-5x5 -n=-25 Both the minues sign will cancel and it will be n=25

OpenStudy (anonymous):

You basically multiply both sides by 5. And the negative signs cancel each other out!

OpenStudy (anonymous):

haha kk thanks peeps

OpenStudy (anonymous):

No probs

OpenStudy (anonymous):

:)